### v2

@@ Appendix B Proofs of Theorems
where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, $\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$, $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and $\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
Proof of Theorem 4.2.
The proof is similar to that of Theorem 5.2 and omitted. ∎
Proof of Lemma 5.1.
Define $r_{\lambda}=(r_{1\lambda},...,r_{n\lambda})^{\prime}$. First, note that
@@
because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 16 and 17. Next,
@@
by Assumptions 7 and 15.
Combining Assumption 17, (B.4), (B.5) and (B.6), we obtain
@@
Now, by Assumptions 15 and 16, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$. Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, writing $\pi_{M}=1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})$ and also using (B.7),
@@
where the Hölder inequality is used to bound $a_{1j,rs},j\geq 4$, as in the proof of Theorem 2 of Kelejian and Prucha (2007). Therefore, we need $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}(1)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}(1)$ as $n\rightarrow\infty$. By Assumption 9(a), we have $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}\left(n^{\eta-1}d_{\gamma}^{2}\right)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}\left(n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\right)$. We then have $a_{1,rs}=o_{p}(1)$ under the sufficient rate in (5.7) as $n\rightarrow\infty$. $a_{2,rs}=o_{p}(1)$ and $a_{3,rs}=o_{p}(1)$ are established exactly as the negligibility of $b_{rs,n}$ and $c_{rs,n}$ is established in the proof of Theorem 1 of Kelejian and Prucha (2007). This concludes the proof. ∎
Proof of Theorem 5.1.
Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$. Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big\|r_{\lambda\delta}\big\|^{2}=o_{p}(1)$ by Assumption 16. Under $H_{0}:\vartheta=0$,
@@
where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, $\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$, $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and $\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$. Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big\|\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 15 and 17. Thus, $\big\|\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big\|\leq\big\|\widehat{\mathcal{U}}_{3}^{-1}\big\|\big\|\mathcal{U}_{3}-\widehat{\mathcal{U}}_{3}\big\|\big\|\mathcal{U}_{3}^{-1}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 6 and 17, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$. Similarly, we have $\big\|\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big\|=o_{p}(1)$, and $\big\|\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big\|=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 7 and 15.
@@
Similar to (B.13) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\vartheta}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 7 and 15, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\vartheta}}}|A_{22}|=o_{p}(1)$. Therefore, we obtain the sufficient rate from (5.7) as $n\rightarrow\infty$ for the required quantities to be negligible. ∎
Proof of Theorem 5.2.
(i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1. We now note that the entries of $\mathscr{S}$ are $s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so $\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i,j$, by Assumptions 15 and 17, and arguments used in the proof of Theorem 5.1 above. Similarly, we also observe that $\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i$. Thus Assumption A(b) is satisfied. Assumption A is satisfied to due to Assumption 17. The conclusion now follows by Theorem A.1.
@@
From Assumptions 15 and 17, we have $\left\|\widehat{\mathscr{D}}_{4}-\mathscr{D}_{4}\right\|=o_{p}(1)$ and

### v3

@@ Appendix B Proofs of Theorems
where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, $\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$ with $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and $\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. $\blacksquare$
(Proof).[Proof of Theorem 4.2] The proof is similar to that of Theorem 5.2 and omitted. $\blacksquare$
(Proof).[Proof of Lemma 5.1] Define $r_{\lambda}=(r_{1\lambda},...,r_{n\lambda})^{\prime}$. First, note that
@@
because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 5.2 and 5.3.
Next,
@@
by Assumptions 3.2 and 5.1.
Combining Assumption 5.3, (B.4), (B.5) and (B.6), we obtain
@@
Now, by Assumptions 5.1 and 5.2, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$. Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, writing $\pi_{M}=1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})$ and also using (B.7),
@@
where the Hölder inequality is used to bound $a_{1j,rs},j\geq 4$, as in the proof of Theorem 2 of Kelejian and Prucha (2007). Therefore, we need $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}(1)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}(1)$ as $n\rightarrow\infty$. By Assumption 3.4(a), we have $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}\left(n^{\eta-1}d_{\gamma}^{2}\right)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}\left(n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\right)$. We then have $a_{1,rs}=o_{p}(1)$ under the sufficient rate in (5.7) as $n\rightarrow\infty$. $a_{2,rs}=o_{p}(1)$ and $a_{3,rs}=o_{p}(1)$ are established exactly as the negligibility of $b_{rs,n}$ and $c_{rs,n}$ is established in the proof of Theorem 1 of Kelejian and Prucha (2007). This concludes the proof. $\blacksquare$
(Proof).[Proof of Theorem 5.1] Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$. Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big\|r_{\lambda\delta}\big\|^{2}=o(1)$ by Assumption 5.2. Under $H_{0}:\vartheta=0$,
@@
where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, $\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$, $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and $\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$. Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big\|\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 5.1 and 5.3. Thus, $\big\|\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big\|\leq\big\|\widehat{\mathcal{U}}_{3}^{-1}\big\|\big\|\mathcal{U}_{3}-\widehat{\mathcal{U}}_{3}\big\|\big\|\mathcal{U}_{3}^{-1}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 3.1 and 5.3, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$. Similarly, we have $\big\|\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big\|=o_{p}(1)$, and $\big\|\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big\|=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 3.2 and 5.1.
@@
Similar to (B.13) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\vartheta}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 3.2 and 5.1, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\vartheta}}}|A_{22}|=o_{p}(1)$. Therefore, we obtain the sufficient rate from (5.7) as $n\rightarrow\infty$ for the required quantities to be negligible. $\blacksquare$
(Proof).[Proof of Theorem 5.2] (i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1. We now note that the entries of $\mathscr{S}$ are $s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so $\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i,j$, by Assumptions 5.1 and 5.3, and arguments used in the proof of Theorem 5.1 above. Similarly, we also observe that $\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i$. Thus Assumption A(b) is satisfied. Assumption A is satisfied to due to Assumption 5.3. The conclusion now follows by Theorem A.1.
@@
From Assumptions 5.1 and 5.3, we have $\left\|\widehat{\mathscr{D}}_{4}-\mathscr{D}_{4}\right\|=o_{p}(1)$ and